> >SOHO DAILY REPORT > >Operations Day - 202/2000 (20 July 00) >Mission Day 1693 > >SOHO FLIGHT OPERATIONS TEAM, NASA/GSFC >Ref. #: SH-FOT-DR-000720 >===================================================================== >PASS #1 Madrid DSS54 0752-1720 >PASS #2 Goldstone DSS27 1655-1942 >PASS #3 Goldstone DSS16 1919-203/0129 >********************************************* > > >PASS #1 Madrid DSS54 0752-1720 > >Anomalies: > >None > >NRT: > >1509-1640 > >Activities Summary: > >- SSR Dump >- Collected MDI-M (0957-1159) >- Collected MDI-H (1159-1645, 1652-1720) >- Collected Ranging Data >- Uplinked delayed load VIRA00721 >- HGA Table Upload >- Hx Limit Update >- Reaction Wheel Speed Monitoring Daily Adjustment >- CRS Drift Estimation Update >- Executed OCD# 1397 - Enable Experiment Offpointing for LASCO >- SSR SEF Reset >------------------ > > >PASS #2 Goldstone DSS27 1655-1942 > >Anomalies: > >G00-0311 1703 UPS went down during maintenance disrupting power to the >operational > string, IRTS and the EOF. FOT failed over to a backup >string. EOF > was down for approximately 32 minutes. > >NRT: > >None > >Activities Summary: > >- Collected MDI-M (1718-1820) >- Collected MDI-H (1655-1718, 1820-1942) >------------------ > > >PASS #3 Goldstone DSS16 1919-203/0129 > >Anomalies: > >G00-0312 1925 VDA failure in D16's front end resulted in the inability >to command > for 31 minutes, DR# G06181 > >G00-0313 2250 Lost uplink due to PA failure at D16. Lost ability to >command for 4 > minutes, DR# G06182 > >NRT: > >2000-2249 >2255-2311 > >Activities Summary: > >- Collected MDI-H (1919-0129) >------------------ > > >-- End of Report -- >